OpenStudy (anonymous):

The number of atoms in 10.0 gram of CaCO3 (100.0g/mol) is?

OpenStudy (anonymous):

n=m/M = 10 g / 100 gmol-1 = 0,1 mol n=N/L --> N= n+L N= 0,1 mol * 6,022*10^23mol -1 = 6,022*10^22 molecules

OpenStudy (anonymous):

See I messed up on the first step. I used 10.0g and multiply that by 1molCaCO3/mass of CaCO3

OpenStudy (anonymous):

you can also use simple equations and rearange them cause that way you make your life easier: n1=m/M n2= N/L n1=n2 m/M=N/L mL=NM N=mL/M